OpenStudy (anonymous):

help please. I have 30 min. to finish my homework and my mom is going to yell at me :( please help . i am begging you

OpenStudy (anonymous):

Choose the equation that represents the graph below: Graph of a line passing through points negative 3 comma 0 and 0 comma 3 y = x - 3 y = -x + 3 y = -x - 3 y = x + 3

OpenStudy (anonymous):

@Jhannybean @nincompoop @pooja195 @ganeshie8

OpenStudy (anonymous):

are the things listed answers?

OpenStudy (anonymous):

@Directrix @@mathstudent55 @mathmath333 @Luigi0210

OpenStudy (anonymous):

what? @idontnowagoodusername just help !

OpenStudy (anonymous):

i'm not giving you the answer but i'll help you through it

OpenStudy (anonymous):

ok well can you :(

OpenStudy (anonymous):

i think it is a or b but i do not know

OpenStudy (anonymous):

okay since your on a time thing is B